Fill in the blank with the appropriate direction (left, right, up, or down).
(a) The graph of y = f(x) + 2 is obtained from the graph of y = f(x) by shifting [????] 2 units.
(b) The graph of y = f(x + 2) is obtained from the graph of y = f(x) by shifting {????} 2 units.

Fill in the blank with the appropriate direction (left, right, up, or down).
(a) The graph of y = f(x) − 9 is obtained from the graph of y = f(x) by shifting [????] 9 units.
(b) The graph of y = f(x − 9) is obtained from the graph of y = f(x) by shifting [????] 9 units.

Fill in the blank with the appropriate axis (x-axis or y-axis).
(a) The graph of y = −f(x) is obtained from the graph of y = f(x) by reflecting in the [????].
(b) The graph of y = f(−x) is obtained from the graph of y = f(x) by reflecting in the [????].

How do I solve these?
Thank you

Answer by Alan3354(69443)

Step 1, stop using all the ?'s
We're not impressed by ????
-------------
For example,
Fill in the blank with the appropriate direction (left, right, up, or down).
(a) The graph of y = f(x) + 2 is obtained from the graph of y = f(x) by shifting [????] 2 units.
Try
(a) The graph of y = f(x) + 2 is obtained from the graph of y = f(x) by shifting ? 2 units. 2 units in what direction?